OpenStudy (anonymous):

If you go from 0 m/s to 50m/s in 5 seconds what is the distance you have traveled PLEASE HELP I just need some clarification

OpenStudy (anonymous):

would the answer be 150 meter as each seconds she goes an average of 10m/s faster

OpenStudy (anonymous):

If you're learning kinematics equation you would use (xf - xi) = vit + (1/2)at^2 We know what our initial velocity is = 0m/s we know what our final veloctiy is =50m/s we know what time it takes = 5s we need to find acceleration. acceleration can be found by (vf - vi) / t

OpenStudy (anonymous):

with all the variables known except for x, you can now find the distance

OpenStudy (anonymous):

THANK YOU

OpenStudy (anonymous):

WAIT what does vit stand for

OpenStudy (anonymous):

intial velocity * time?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

thanks for the clarification

OpenStudy (anonymous):

np

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

you a life saver

OpenStudy (anonymous):

would the answer be 2.5 or did i mess up

OpenStudy (anonymous):

oh shoot i messed up

OpenStudy (anonymous):

i haven't done the calculated haha.

OpenStudy (anonymous):

wait so if x equals the distance traveled.........

OpenStudy (anonymous):

yeah, you know that initial distance is 0m so you're looking for the final distance

OpenStudy (anonymous):

wait so i got x=2.5secs

OpenStudy (anonymous):

(x - 0) = 0m/s ( 5seconds) + 1/2 (10m/s^2) (5seconds) ^2

OpenStudy (anonymous):

and x = displacement

OpenStudy (anonymous):

i still got x=2.5 secs

OpenStudy (anonymous):

are the seconds supposed to cancel out?

OpenStudy (anonymous):

AHHH

OpenStudy (anonymous):

you just plug it in lol

OpenStudy (anonymous):

what are you stuck on?

OpenStudy (anonymous):

i think i am simplifying wrong

OpenStudy (anonymous):

you don't need to simplify anything hahahaha

OpenStudy (anonymous):

I don't?

OpenStudy (anonymous):

write down the given for me

OpenStudy (anonymous):

5 seconds is the given right?

OpenStudy (anonymous):

yea, what else

OpenStudy (anonymous):

the velocity

OpenStudy (anonymous):

you also have one more

OpenStudy (anonymous):

its kinda hidden, but your initial distance is 0m

OpenStudy (anonymous):

acceleration?

OpenStudy (anonymous):

oh yeah

OpenStudy (anonymous):

okay so you have t=5s vi = 0m/s vf = 50m/s xi = 0m

OpenStudy (anonymous):

you're using this kinematic equation (xf - xi) = vit + 1/2at^2

OpenStudy (anonymous):

what are you missing?

OpenStudy (anonymous):

xf?

OpenStudy (anonymous):

and?

OpenStudy (anonymous):

vf?

OpenStudy (anonymous):

acceleration, but you can find accerlation with vf - vi / t

OpenStudy (anonymous):

and a = 10m/s^2 so now all you have left is xf

OpenStudy (anonymous):

(xf - 0 ) = (0) + 1/2 (10) (5) ^2

OpenStudy (anonymous):

is it 125!

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

YESSSSSSSSS I DID IT THANK YOU

OpenStudy (anonymous):

what grade are you in

OpenStudy (anonymous):

i think i misinterpreted t^2 as seconds squared

OpenStudy (anonymous):

I am in 9th

OpenStudy (anonymous):

i know its sad

OpenStudy (anonymous):

did you guys learn kinematic equation yet?

OpenStudy (anonymous):

Its a little sad I got messed up by seconds squared

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how does your teacher go about doing this question

OpenStudy (anonymous):

Sasha drives her hovercraft from 0m/s to 50 m/s in 5 sec, how far does she go

OpenStudy (anonymous):

thats the orginal question

OpenStudy (anonymous):

how does your teacher solve it?

OpenStudy (anonymous):

its hw we havent gone over it yet

OpenStudy (anonymous):

surprise your teacher!!!! lol

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you have the patience of a sain

OpenStudy (anonymous):

sorry i meant patience of a saint

OpenStudy (anonymous):

thanks I owe you

OpenStudy (anonymous):

study hard!!! don't slack

OpenStudy (anonymous):

Got it! Good luck in whatever you study